OpenStudy (xxhelpmehplzxx):

Will give medal. 3𝑥 – 57 + 𝑥 = 63 Solve for x.

OpenStudy (abdullahm):

First add like terms What is 3x + x = ?

OpenStudy (xxhelpmehplzxx):

4x?

Mehek (mehek14):

Correct 4x - 57 = 63 We need to get rid of - 57 so add 63 + 57 = ?

OpenStudy (xxhelpmehplzxx):

120

Mehek (mehek14):

4x = 120 divide by 4 120/4 = ?

OpenStudy (xxhelpmehplzxx):

30.

Mehek (mehek14):

x = 30 :)

OpenStudy (xxhelpmehplzxx):

thank you

Mehek (mehek14):

yw :)